OpenStudy (anonymous):

The volume of a cylinder is given by the formula V=pi r^2h. What is the volume of the cylinder below. 12cm going up and 9 cm on the bottom

OpenStudy (anonymous):

s=23

OpenStudy (anonymous):

were do u get s from ?

OpenStudy (anonymous):

V=pi*(r^2)*h V=pi*((9/2)^2)*12 V=pi*(4.5^2)*12 V=763.407 (or 763.41) cm^3

OpenStudy (anonymous):

mvm I guess i accidently replied to the wrong problem...